Two kg of a substance undergoes the specified process in a Cylinder-piston arrangement:

p1 = 6 bar; V1 = 0.2 m³; p2 = 2 bar; V2 = 0.6 m³,

Determine the work done (W) in each case.

(a) p varies as a linear function of V. (b) pV = a constant. (C) p remains constant till the volume reaches V3 = 0.3m³; pV^n = a Constant after that4 (You have to determine the value of n.)
1
Expert's answer
2020-10-01T10:07:57-0400

solution

P1= "6\times10^5N/m^2" V1 "=0.2m^3"

P2="2\times10^5N/m^2" v2 "=0.6m^3"

(a)if p varies linearly as function of V then


"P=a+bV"


putting the value of P1 ,V1 and P2 V2 in above equation


"6\times10^5=a+b\cdot 0.2\\2\times10^5=a+b\cdot 0.6"


from above equation a="8\times10^5"

b="-10\times10^5"

so work done by gas can be calculated as

"W=\int pdv"

"W=\int_{0.2}^{0.6}(8\times10^5-10^6V)dV"


"W=1.6\times10^5J"


(b)if PV=c then work done can be given as


"W=P_2V_2ln\frac{V_2}{V_1}"


"W=2\times10^5\times0.6ln\frac{0.6}{0.2}"


"W=0.56\times10^5J"

(C)if pvn=constant then


"P_1V_!^n=P_2V_2^n"


by putting the value of Paressure and volume

"6\times10^5\times0.3^n=2\times10^5\times0.6^n\\n=1.58"


when n=1.58 then workdone is given as

"W=\frac{P_2V_2-P_1V_1}{1-n}"


"W=\frac{(2\times10^5\times0.6)-(6\times10^5\times0.3)}{1-1.58}"


"W=1.03\times10^5J"


so work done during volume expantion from 0.3m^3 to 0.6m^3 is "1.03\times10^5J ."